Q3. Suppose that people consume only three goods, as shown in this table: Tennis Balls Golf Balls 100 Plus 2006 Price RM2.00 RM4.00 RM1.00 2006 Quantity 100 100 200 2007 Price RM2.00 RM6.00 RM2.00 2007 Quantity 100 100 200 a. What is the percentage change in the price of each of the three goods? b. Using a method that similar to the consumer price index, compute the percentage change in the overall price level. c. If you were to learn that a bottle of 100 Plus increased in size from 2006 and 2007, should that information affect your calculation of the inflation rate? If so, how? d. If you were to learn that a bottle of 100 Plus introduced new flavors in 2007, should that information affect your calculation of the inflation rate? If so, how?


1
Expert's answer
2021-11-22T10:01:36-0500

a.

b.

The cost of the market basket in 2014 is "(\$2 × 100) + (\$4 × 100) + (\$1 × 200) = \$200 + \$400 + \$200 = \$800." The cost of the market basket in 2015 is "(\$2 × 100) + (\$6 × 100) + (\$2 × 200) = \$200 + \$600 + \$400 = \$1,200." The percentage change in the cost of the market basket from 2014 to 2015 is "\frac{(1,200 – 800)}{800} × 100\% = 50\%."

c. Knowing that the size of a bottle of Gatorade rose from 2014 to 2015 would cause you to lower your inflation rate estimate. Because a bottle of Gatorade is suddenly worth more than it was previously. On a per-ounce basis, the comparison should be made.


b. Knowing that Gatorade introduced additional flavors in 2015 would lower your inflation rate estimate. Because more flavors make people feel better. As a result, this would be regarded as a change in quality.